OpenStudy (anonymous):

Convert into slope intercept form: 3x-y=1

OpenStudy (solomonzelman):

y=mx+b is the slope intercept form 3x -y = 1 -3x -3x -y= 1 - 3x that's same as -y= - 3x +1 divide everything by -1, and you get y=3x-1